OpenStudy (nbarrera):

Factor the following expression completely. x4-81

OpenStudy (bloomlocke367):

okay, do you know how to factor? I can help you

OpenStudy (nbarrera):

\[x^{4}-81\]

OpenStudy (bloomlocke367):

Yes, but do you know how to factor?

OpenStudy (nbarrera):

no

OpenStudy (bloomlocke367):

what if that just said\[x^2-81\] would you know how to do it then?

OpenStudy (nbarrera):

it would be 9?

OpenStudy (nbarrera):

idk

OpenStudy (bloomlocke367):

okay, that's a good start.

OpenStudy (bloomlocke367):

So, you need to treat that fourth degree polynomial as just a quadratic.

OpenStudy (nbarrera):

Then what

OpenStudy (bloomlocke367):

We can factor a difference of fourth powers (and higher powers) by treating each term as the square of another base, using the power to a power rule. For example, to factor x 4 - y 4 , we treat x 4 as (x 2)2 and y 4 as (y 2)2 . Thus, x 4 - y 4 = (x 2)2 - (y 2)2 = (x 2 + y 2)(x 2 - y 2) = (x 2 + y 2)(x + y)(x - y) .

OpenStudy (bloomlocke367):

does that make sense? or do I need to explain further?

OpenStudy (nbarrera):

Yes thx.

OpenStudy (bloomlocke367):

let me know what you get in the end. I'll tell you if you're correct

OpenStudy (bloomlocke367):

To begin with, you should have\[(x^2+9)(x^2-9)\]

OpenStudy (bloomlocke367):

@nbarrera